### v1

# Optimal Parallel $\lambda$-Reduction

Abstract
I present a model of universal parallel computation called $\Delta$-Nets, and a method to translate $\lambda$-terms into $\Delta$-nets and back. Together, the model and the method constitute an algorithm for optimal parallel $\lambda$-reduction, solving the longstanding enigma with groundbreaking clarity. I show that the $\lambda$-calculus can be understood as a projection of $\Delta$-Nets—one that severely restricts the structure of sharing, among other drawbacks. Unhindered by these restrictions, the $\Delta$-Nets model opens the door to new highly parallel programming language implementations and computer architectures that are more efficient and performant than previously possible.
Interactive Demo and Source Code
1.  Introduction
The $\lambda$-calculi are beautifully simple yet powerful models of universal computation. Consisting of abstractions, variables, and applications, $\lambda$-terms can express any computable function [Chu36, Tur36, Tur37, Chu41]. In addition to being central pillars in computation theory, the $\lambda$-calculi also constitute practical frameworks underpinning all functional programming languages. Four $\lambda$-calculi are pertinent to this paper—the three substructure $\lambda$-calculi [Jac93], and the full $\lambda$-calculus [Bar84]:
  • $\bullet$
    $\lambda L$-calculus: the linear $\lambda$-calculus, in which every bound variable occurs exactly once.
  • $\bullet$
    $\lambda A$-calculus: the affine $\lambda$-calculus, in which every bound variable occurs either once or not at all.
  • $\bullet$
    $\lambda I$-calculus111the original $\lambda$-calculus defined by Church [Chu36, Chu41].: the relevant $\lambda$-calculus, in which every bound variable occurs at least once.
  • $\bullet$
    $\lambda K$-calculus: the full $\lambda$-calculus in which bound variables can occur any number of times.
The $\lambda L$-calculus can be regarded as a cornerstone which can be extended in one of three ways: with erasure (analogous to weakening in logic), resulting in the $\lambda A$-calculus; with sharing (analogous to contraction in logic), resulting in the $\lambda I$-calculus; or with both erasure and sharing, resulting in the full $\lambda K$-calculus. Additionally, the $\lambda K$-calculus can also be obtained by extending the $\lambda A$-calculus with sharing, or the $\lambda I$-calculus with erasure. These relationships are illustrated in Figure 1.
$\lambda L$$\lambda A$$\lambda I$$\lambda K$erasuresharing
Figure 1: The relationships between the four $\lambda$-calculi in terms of erasure and sharing.
In the $\lambda$-calculi, a reduction strategy is optimal if and only if it reaches the normal form (if it exists) without perfoming any unnecessary reduction steps [Lev78, Lev80]. There are two types of unnecessary reductions:
1. Reduction of a later-discarded subexpression.
2. Reduction of a duplicated subexpression.
The first type can only occur with erasure, i.e., when some abstractions do not make use of their bound variables. Whenever such an abstraction is applied, the argument is discarded, and any reductions previously performed in the argument subexpression are rendered unnecessary. Naturally, this cannot happen in the $\lambda L$-calculus nor in the $\lambda I$-calculus. This first type of unnecessary reduction can be entirely avoided in the $\lambda A$-calculus and in the $\lambda K$-calculus by adhering to some reduction orders, including normal order reduction [Bar87].
The second type of unnecessary reduction can only occur with sharing, i.e., when some abstractions’ bound variables occur multiple times. Whenever such an abstraction is applied, the argument is duplicated, and any reducible expressions in the argument subexpression are duplicated with it. Naturally, this cannot happen in the $\lambda L$-calculus nor in the $\lambda A$-calculus. Lévy has shown that there are $\lambda$-terms with sharing for which no reduction order is optimal [Lev78, Lev80]. The $\lambda$-calculus, as a sequential substitution machine, is therefore inadequate to express optimal reduction for all $\lambda$-terms. Does a more fundamental model of computation exist which is able to express optimal reduction for all $\lambda$-terms?
In order to avoid sharing-related unnecessary reductions in the $\lambda I$- and $\lambda K$-calculi, it is useful to represent $\lambda$-terms as graphs. Identical subexpressions can then be represented by the same shared subgraph, which only needs to be reduced once. The process of reduction is then expressed as a sequence of graph operations—a technique known as graph reduction [Wad71]. In the $\lambda$-calculi, applying a function destructively modifies its body. This presents a challenge in graph reduction when a function is shared, since it may be applied any number of times, each with a different argument. A simple solution is to duplicate the shared function’s entire subgraph before applying it. This solution, however, leads to the second type of unnecessary reduction because it duplicates reducible expressions. It’s possible to mitigate the number of duplicated reducible expressions, and thus of unnecessary reductions, by sharing, instead of duplicating, the function’s maximal free subgraphs—the largest subgraphs in the function’s body that don’t make use of the function’s bound variable [Wad71]. An ordering can also be imposed such that all reducible expressions inside a function are first reduced, and only then is the function duplicated (while sharing its maximal free subgraphs). However, a critical problem remains: this procedure never terminates in cyclic graphs, and cyclic graphs can arise from non-cyclic ones through regular reduction [Wad71]. In [Wad71] this is resolved by duplicating subgraphs whenever necessary to avoid cycles.
The first algorithms for $\lambda$-calculi reduction which don’t perform any unnecessary $\beta$-reductions were proposed in [Lam89] and [Kat90]. The common core idea introduced was that of interior sharing of subgraphs. Interior sharing enables the incremental duplication of shared functions, which, in turn, fully prevents sharing-related unnecessary $\beta$-reductions. The challenge of interior sharing lies in the management, throughout reduction, of multiple simultaneous sharing contexts, each of which can fully or partially overlap any number of others. Moreover, sharing contexts can be recursive—a shared term can be referenced from inside itself any number of times. In [Lam89], interior sharing is accomplished through the use of explicit fan-in and fan-out nodes. Remarkably, the complex challenge then boils down to a simple question: when a particular fan-in meets a particular fan-out, should they annihilate one another or duplicate one another? In [Lam89], this is solved by associating a non-negative integer with each fan and introducing three delimiter node types to regulate fan numbers through graph-rewriting rules. When two fans meet, they are annihilated if and only if their numbers are equal, and they duplicate one another otherwise (Figure 2). The delimiters realize a notion of enclosure around fans, ultimately ensuring that only fans belonging to the same enclosure annihilate one another.
$a=b$$a$$b$$a\neq b$$b$$b$$a$$a$
Figure 2: Interacting fans annihilate one another if they are equal, otherwise they duplicate one another.
Concurrently, a graphical meta-model of parallel computation called interaction nets was introduced in [Laf90]. It was later refined and expanded in [Laf97], once its significance was understood better. Models adhering to this framework are called interaction systems. An interaction system is specified by a set of agent types and a set of interaction rules—local graph rewriting rules that operate on pairs of connected agents. Agents are nodes that have one principal port and any number of auxiliary ports, including zero. When two agents are connected via their principal ports they form an active pair, and an interaction rule can then be applied to that pair. Since each agent can only be part of a single active pair at a time, interaction systems possess a one-step diamond property, which I denote as perfect confluence [^2] Some authors denote this as strong confluence, but strong confluence has historically denoted a weaker property which is not one-step. I propose using the term perfect confluence for the one-step variant to avoid confusion.. An extraordinary consequence of perfect confluence is that every normalizing interaction order produces the same result in the same number of interactions. Additionally, since interaction rules are local, they can be applied simultaneously without synchronization. These properties make interaction systems highly suitable for expressing optimal parallel algorithms.
An interaction system for $\lambda$-calculi reduction based on [Lam89], along with a translation method from and to $\lambda$-terms, was proposed in [GAL92a, GAL92b]. A similar interaction system, combined with a different translation method, was presented in [AG98]. These interaction systems employ indexed agents called brackets and croissants, which increment and decrement, respectively, the index of higher-index agents they interact with. As explained in detail in [AG98], there are many ways to translate $\lambda$-terms into interaction nets using brackets and croissants. These agents effectively delimit sharing scopes, and, as such, are also referred to as delimiters. A comprehensive analysis and comparison of the computational efficiency of these algorithms, including the original [Lam89], was supplied in [LM99]. Unfortunately, all these algorithms have frustrating characteristics that profoundly undermine reduction performance. Critically, delimiters accumulate during reduction until delimiter interactions completely overwhelm fan interactions. Additionally, delimiters are often present in nets associated with $\lambda$-terms that have no sharing, serving no purpose there.
A more recent interaction system for $\lambda$-calculi reduction, called Lambdascope, was presented in [OL04]. It uses five agent types: an abstractor, an applicator, a duplicator (an indexed fan), an eraser, and an indexed delimiter. In Lambdascope, applying an abstraction effectively spawns two zero-indexed delimiters. This process, along with other interaction rules, preserves the scope of each abstraction after it has been applied. This scope invariant, in turn, ensures interacting duplicators annihilate and commute when appropriate. Delimiters move outward, expanding each scope as much as possible. Sibling scopes eventually coalesce, which helps reduce the number of total scopes and thus the number of total delimiters. However, siblingless scopes and their delimiters are preserved perpetually. As a result, Lambdascope also suffers from a devastating accumulation of delimiters. Additionally, since there’s no mechanism to decrement indexes, they grow without bound as reduction progresses.
For example, when reducing the non-normalizing $\lambda$-term $(\lambda x.x\,x)(\lambda y.y\,y)$ with these algorithms, every iteration creates new delimiters. Each new delimiter takes up additional memory (space), and leads to additional unnecessary interactions (time). Additionally, the existing algorithms fail to establish a global partial ordering, which is absolutely critical to ensure that all nets associated with normalizing $\lambda$-terms normalize. Not only are the existing algorithms unsatisfactory from a theoretical perspective, their inefficiencies preclude them from being used at the core of programming language implementations.
In this paper, I present a model of universal parallel computation called $\Delta$-Nets, and a method to translate $\lambda$-terms into $\Delta$-nets and back. Together, the system and translation method constitute an algorithm for optimal parallel $\lambda$-reduction. As an interaction system, the core of the model is perfectly confluent: every normalizing reduction order produces the same result in the same number of steps. The core model is then necessarily extended with non-interaction canonicalization rules, which depart from the interaction paradigm. In order to minimize total number of reductions and guarantee normalization for all nets associated with normalizing $\lambda$-terms, a global partial reduction ordering needs to be followed, and is specified. The $\Delta$-Nets algorithm solves the longstanding enigma of optimal $\lambda$-calculi reduction with groundbreaking clarity. Instead of making use of delimiters, sharing is expressed through a single agent type which allows any number of auxiliary ports, called a replicator. Each instance of a replicator incorporates information that in previous models was spread across multiple agents, such as indexed fans and delimiters. This consolidation of information enables simplifications that were previously unfeasible, and leads to constant memory usage in the reduction of $(\lambda x.x\,x)(\lambda y.y\,y)$, for example. Finally, I show that the $\lambda$-calculus can be understood as a projection of $\Delta$-Nets. The additional degrees of freedom in $\Delta$-Nets allow it to realize optimal reduction in the manner envisioned by Lévy, i.e., no reduction operation is performed which is rendered unneccessary later, and no reduction operation which is necessary is performed more than once.
2.  Core Interaction System
At its core, a $\Delta$-net is an interaction net with three agent types: a fan, an eraser, and a replicator.
[…]
Figure 3: The three $\Delta$-Nets agent types.
Fans have two auxiliary ports, and erasers have none. Each replicator can have a different natural number of auxiliary ports, and each of those ports has an associated integer called a level delta. Additionally, each replicator has an associated non-negative integer called a level. As a point of reference, an agent’s auxiliary ports are ordered in a clockwise orientation, i.e., from left to right when the principal port points down.
[…]
Figure 4: The core $\Delta$-Nets interaction rules.
Fans are sufficient to express optimal parallel reduction in the $\lambda L$-calculus. Erasers are needed to express optimal parallel reduction involving erasure. Replicators are needed to express optimal parallel reduction involving sharing. As such, the $\Delta$-Nets core interaction system decomposes perfectly into three overlapping subsystems, each analogous to a substructure $\lambda$-calculus. The full $\Delta$-Nets system may also be referred to as $\Delta K$-Nets. The four systems are:
  • $\bullet$
    $\Delta L$-Nets: the linear $\Delta$-Nets subsystem, which only uses fans, and expresses optimal parallel reduction in the $\lambda L$-calculus.
  • $\bullet$
    $\Delta A$-Nets: the affine $\Delta$-Nets subsystem, which uses fans and erasers, and expresses optimal parallel reduction in the $\lambda A$-calculus.
  • $\bullet$
    $\Delta I$-Nets: the relevant $\Delta$-Nets subsystem, which uses fans and replicators, and expresses optimal parallel reduction in the $\lambda I$-calculus.
  • $\bullet$
    $\Delta K$-Nets: the full $\Delta$-Nets system, which uses fans, erasers and replicators, and expresses optimal parallel reduction in the $\lambda K$-calculus.
When equal agents interact, they annihilate one another. In general, two replicators are equal if and only if they have the same level, number of auxiliary ports, and level deltas. However, when a $\Delta$-net is constructed from a $\lambda$-term via the translation method presented in the next section, interacting replicators that have the same level are guaranteed to be equal. As such, only replicator levels need to be compared for equality in practice. Equal-agent interactions are called annihilations.
When distinct agents interact, each agent travels through and past the other and is potentially copied or erased in the process, depending on how many auxiliary ports the other agent has. Since an eraser has no auxiliary ports, it erases every agent it interacts with. As such, distinct-agent interactions involving erasers are called erasures. The remainder of distinct-agent interactions are called commutations. When a replicator interacts with a fan, the replicator travels through and out of the fan’s two auxiliary ports, resulting in two exact copies of the replicator. Simultaneously, the fan travels through and out of all of the replicator’s auxiliary ports, resulting in a fan for each replicator auxiliary port.
When two distinct replicators interact, the lower-level one replicates the higher-level one once for each of the lower-level one’s auxiliary ports, while, simultaneously, the higher-level one duplicates the lower-level one once for each of the higher-level one’s auxiliary ports. Note that whereas a duplication produces exact copies, a replication produces copies that may or may not be exact. Replicators are so named because the copies they produce of other replicators are, by design, not necessarily exact. Each resulting replica of the higher-level replicator may have a different level, determined by the level delta associated to the auxiliary port of the lower-level replicator that it travels out of. The level of each resulting replica is the sum of the level of the original higher-level replicator and the appropriate level delta of the lower-level replicator. The $\Delta$-Nets interaction rules are illustrated in Figure 4.
3.  From $\lambda$-terms to $\Delta$-nets
Definition.
Let $\Sigma=\{L,A,I,K\}$. For all $S\in\Sigma$, let $\Lambda_{S}$ be the set of all $\lambda S$-terms and $\Delta_{S}$ be the set of all $\Delta S$-nets. For all $S\in\Sigma$, there exists a bijection $\phi_{S}:\Lambda_{S}\rightarrow\Delta_{S}^{c}$ which maps every $\lambda S$-term to a canonical $\Delta S$-net.
$\Delta_{S}^{c}$ =$=\{\phi_{S}(\lambda_{S})\mid\forall\lambda_{S}\in\Lambda_{S}\}$ (canonical $\Delta S$-nets)
$\Delta_{S}^{p}$ =$=\{\delta_{S}^{p}\mid\forall\delta_{S}^{c}\in\Delta_{S}^{c},\ \delta_{S}^{c}\stackrel{{\scriptstyle\Delta^{*}}}{{\to}}\delta_{S}^{p}\}$ (proper $\Delta S$-nets)
$\Delta_{S}^{c}\subseteq\Delta_{S}^{p}\subset\Delta_{S}$
The bijections $\phi_{S},\ \forall S\in\Sigma$, are defined inductively, with the rules for the general $\phi_{K}$ case illustrated in Figure 5. In the description that follows, the minor differences in $\phi_{L}$, $\phi_{A}$, and $\phi_{I}$ are noted where appropriate. Each outer dashed rectangle in Figure 5 contains a $\Delta$-net fragment that represents the $\lambda$-term specified above it. Each inner dashed rectangle is a slot for the $\Delta$-net that represents the inner term. Each thick vertical dashed line represents a non-negative integer number of parallel wires (including zero). Every dashed rectangle, outer or inner, has the same interface—a single wire entering at the top and a non-negative integer number of wires leaving at the bottom. Each $\lambda$-term has a subscript associated with it, which represents the level of that term. The level of the outermost term is set to zero, which inductively sets all other levels. The level of an application’s argument is one greater than that of the application itself, and the level of a replicator is one greater than that of its associated abstraction. These levels are ultimately used to determine the level and level deltas of replicators.
$[\lambda x.M]_{l},\ x\not\in FV(M)$$\lambda$$[M]_{l}$$d_{i}=l_{i}-(l+1)$ $[\lambda x.M]_{l},\ x\in FV(M)$$\lambda$$\ldots$$[M]_{l}$$l_{0}$$l_{1}$$l_{n}$$d_{0}$$d_{1}$$d_{n}$$l+1$$[MN]_{l}$@@@@$[M]_{l}$$[N]_{l+1}$$[x]_{l},\ x\not\in FV(*)$$l$$[x]_{l},\ x\in FV(*)$$x$
Figure 5: The inductive definition of $\phi_{K}$, a bijection which translates $\lambda K$-terms into canonical $\Delta K$-nets.
There are two $\Delta$-net fragments for variables, one for variables that are free in the outermost $\lambda$-term ($x\in FV(*)$), and one for variables that are not ($x\not\in FV(*)$), i.e., that are bound by some abstraction. The free-variable fragment contains a single-port (non-agent) node, which is represented by the name of the associated free variable in the $\lambda$-term. The bound-variable fragment is just a vertical wire. An instance of the bound-variable fragment which represents the $i$th occurrance of a bound variable in the associated $\lambda$-term has its bottom wire endpoint connected to the $i$th auxiliary port of the replicator that shares that variable. The bottom wire endpoint of each bound-variable fragment has a level associated with it, which is equal to the level of the variable. These wire endpoint levels are referenced in the fragment that features a replicator, in the definition of the level deltas $d_{i}$. The level delta associated to an auxiliary port of a replicator is equal to the level of the wire connected to that auxiliary port minus the level of the replicator.
There are also two $\Delta$-net fragments for abstractions, one for those which use their bound variables and one for those which don’t. In both, the abstraction itself is represented by a fan pointing up, labelled with a $\lambda$ [^3] Fan labels are just a visual aid and do not affect how fans interact with other agents.. The principal port of the fan is the parent port of the abstraction. The first auxiliary port of the fan is a child port, connected to the $\Delta$-net that represents the body of the abstraction. The second auxiliary port represents the variable of the abstraction, and since it is connected to the parent of the variable, it is another parent port. If an abstraction doesn’t use its bound variable, the fan’s second auxiliary port is connected to an eraser. The fragment with this eraser is only needed when translating expressions from $\lambda$-calculi with erasure, and, as such, it is exclusively part of $\phi_{A}$ and $\phi_{K}$. When translating expressions from $\lambda$-calculi with sharing, if an abstraction uses its bound variable, then the fan’s second auxiliary port is generally connected to the principal port of a replicator, which shares the abstraction’s variable among the bound-variable fragments that represent its various occurances in the $\lambda$-term. However, in a canonical $\Delta$-net, a replicator with a single auxiliary port and a level delta of zero, regardless of the replicator’s level, is equivalent to a wire. Therefore such replicators are never created in the first place, and the appropriate bound-variable fragment’s bottom wire is connected directly to the abstraction fan’s second auxiliary port instead. In $\phi_{L}$ and $\phi_{A}$, which translate $\lambda$-terms without sharing, the fragment with the replicator is modified such that the replicator is substituted by a single wire in the same way.
Finally, there is a fragment for applications, in which the application itself is represented by a fan pointing down, labelled with an @@@@3. The first auxiliary port of the fan is the parent port of the application. The principal port of the fan is a child port, connected to the $\Delta$-net that represents the function $\lambda$-term. The second auxiliary port is another child port, connected to the $\Delta$-net that represents the argument $\lambda$-term. In addition to the fragments shown in Figure 5, every canonical $\Delta$-net has a single root (non-agent) node, represented by a small circle, connected to the top of the outermost $\Delta$-net fragment. Together, the root node and the free variable nodes constitute the interface of a canonical $\Delta$-net.
Note that both applications and abstractions are represented by fans, and $\beta$-reduction is expressed through fan annihilation. Whether a specific fan in a $\Delta$-net represents an abstraction or an application can always be determined by inspecting the net (without fan labels) and tracing paths from the root node. While abstraction fans have two parent ports and one child port, application fans have one parent port and two child ports. Every port of any node (including non-agent nodes) in a proper $\Delta$-net is either a parent port or a child port, and every wire connects a parent port with a child port. Although this duality has no direct bearing on the interaction process, it is clearly present, and it is instrumental in some discussions.
All replicators in a canonical $\Delta$-net are unpaired fan-ins: each auxiliary port is a parent port and the principal port is a child port. However, during reduction, fan-out replicators may be produced. In a fan-out replicator, the principal port is a parent port and each auxiliary port is a child port. Every commutation between a fan and a replicator (either a fan-in or a fan-out) always produces a fan-in and a fan-out. While every fan-out is paired with at least one upstream fan-in, the converse is not true: fan-ins may or may not be paired. Locally determining this pairing efficiently is the purpose of the level delta system.
4.  From $\Delta$-nets back to $\lambda$-terms
The only interaction rule in $\Delta L$-Nets is fan annihilation, which expresses $\beta$-reduction. Therefore if $n$ $\beta$-reductions normalize a $\lambda L$-term $t$, then $n$ interactions normalize the $\lambda L$-net $\phi_{L}(t)$. As a result, in the $\Delta L$-Nets system, all proper nets are canonical, and interactions can be applied in any order, with perfect confluence. Compared to $\Delta L$-Nets, the other $\Delta$-Nets systems have additional interaction rules, which don’t have $\lambda$-calculus analogues.
The $\Delta A$-Nets system includes three core interaction rules: fan annihilation, fan erasure, and eraser annihilation. While in the $\lambda$-calculus a single $\beta$-reduction can erase a term of arbitrary size, erasure in $\Delta$-Nets is a computational process whose space and time complexity depends, respectively, on the size and structure of the net being erased. This makes $\Delta$-reductions a more useful measure of computational complexity than $\beta$-reductions.
Applying an abstraction which doesn’t use its bound variable results in an eraser becoming connected to a parent port. Such an eraser can erase abstraction fans and fan-out replicators, but when it reaches the parent port of an application fan, for example, which is an auxiliary port, the erasure process stops. As a result, absent additional rules, irreducible subnets may be produced which don’t have a $\lambda$-calculus analogue. For example, applying an abstraction which doesn’t use its bound variable to an argument which only uses globally-free variables produces a subnet which is disjointed from the root. In order to eliminate all such subnets, two canonicalization rules are introduced: fan decay and free-variable erasure. These rules are illustrated in Figure 6.
$fv$
Figure 6: The $\Delta A$-Nets canonicalization rules:
fan decay and free-variable erasure.
Fan decay allows the erasure of application fans via their parent ports and also of abstraction fans via their body ports. Free-variable erasure allows the erasure of unused free variables. The fan decay rule does not need to differentiate between abstraction and application fans. If an eraser becomes connected to the body port of an abstraction fan, there will be a counterpart eraser somewhere above in the net, connected to a parent port, such that it is safe to decay the abstraction fan.
In the $\Delta A$-Nets system, all three core interaction rules and the two canonicalization rules can be applied in any order, always producing the same result. However, in order to minimize the total number of reduction steps, the fan decay rule should only be applied when the fan annihilation rule cannot be applied. Since in this system only the application of the fan annihilation rule can lead to the creation of new fan-fan active pairs, this partial ordering ensures that all fan annihilations are performed before all fan decays. When no further reduction operation can be applied and the net is normal, it becomes a canonical $\Delta A$-net.
[…]
Figure 7: The aux fan replication canonicalization rule in $\Delta I$- and $\Delta K$-Nets.
During reduction of $\Delta I$- and $\Delta K$-Nets, replicators can combine into replicator trees. A replicator tree is a subnet containing only replicator agents such that each one’s principal port is connected to an auxiliary port of another, except for the tree’s root replicator. In general, replicators that belong to the same tree cannot be merged during reduction because they may be paired. However, if consecutive replicators in a replicator tree are known to be unpaired, they can be safely merged together. Unpaired replicator merging is a canonicalization rule present in both $\Delta I$- and $\Delta K$-Nets. Additionally, in the $\Delta K$-Nets system, it is possible to eliminate unpaired replicators’ auxiliary ports which are connected to erasers (partial unpaired replicator decay). If all auxiliary ports of a replicator are connected to erasers, the entire replicator (regardless of its status) is erased and replaced by an eraser (complete replicator decay). If, as part of partial unpaired replicator decay, an unpaired replicator is left with a single auxiliary port with a level delta of zero, it is replaced by a wire, as it is equivalent to one. These replicator decay rules are $\Delta K$-Nets canonicalization rules.
Replicators start as unpaired, and this status can be propagated across interactions. When an unpaired replicator interacts with a fan, the status changes and the two resulting replicators are no longer unpaired. If an unpaired replicator (A) is connected to a consecutive replicator (B) of unknown status via an auxiliary port, and a certain local constraint is met, then the consecutive replicator can be determined to be unpaired, and the two can then be merged. The constraint is met when the second replicator’s level is greater than or equal to the first replicator’s level, but no greater than the first replicator’s level plus the level delta of the auxiliary port that connects them: $0\leq l_{B}-l_{A}\leq d$. Under this constraint, no replicator is able to interact with the second replicator before the first replicator is annihilated. Since the first replicator is unpaired, it can never be annihilated, and the second one must be unpaired as well.
Meging unpaired replicators always reduces the total number of interactions and the total number of agents, improving space and time efficiency. The optimal partial ordering among replicator commutations is one where active pairs with unpaired replicators are reduced last (specially active pairs of two unpaired replicators). This gives a chance for replicators of unknown status to accumulate behind unpaired replicators and potentially be determined to be unpaired, which would allow them to be merged together. Unpaired replicators can be merged lazily, as part of the fan replication and replicator commutation core interaction rules.
It is beneficial to limit replicators to have at most two auxiliary ports. This imposes the smallest possible upper bound on interaction rule complexity and agent size, making the total number of interactions an effective measure of time complexity, and the total number of agents an effective measure of space complexity. A tree of such replicators can stand in for any replicator with any number of auxiliary ports. There are many ways of building such trees, each with different trade-offs.
The $\Delta I$-Nets system includes all core interaction rules which involve fans and replicators and don’t involve erasers. Additionally, it includes two canonicalization rules: unpaired replicator merging and aux fan replication (illustrated in Figure 7). The $\Delta I$-Nets core interaction rules with lazy unpaired replicator merging, and the aux fan replication canonicalization rule can be applied in any order, always producing the same result. However, in order to minimize the total number of reduction steps, the reduction process needs to be split in two phases. In the first phase the $\Delta I$-Nets core rules with lazy unpaired replicator merging are applied in any order until no active pair is left. The reduction process then switches to the second phase, in which the aux fan replication rule replaces the core fan replication rule. This second phase can be alternatively understood as modifying all fans such that their first auxiliary port becomes the principal port. Similarly to the fan decay rule, the aux fan replication rule does not need to differentiate between abstraction and application fans. If a replicator becomes connected to the body port of an abstraction fan, there will be counterpart replicators positioned such that it is safe to aux-replicate the abstraction fan. This process transforms the sharing structures such that all fan-out replicators are eliminated, all appropriate subnets are replicated, and all fan-in replicators accumulate at the variable port of abstraction fans. Finally, consecutive replicators can be merged since they are all unpaired, resulting in a canonical $\Delta I$-net.
The $\Delta K$-Nets system includes all core interaction rules (augmented with lazy unpaired replicator merging), and all the canonicalization rules that are part of $\Delta A$-Nets and $\Delta I$-Nets. Additionally it includes the replicator decay rules mentioned previously. A complex partial ordering of reduction operations is required in order to minimize the total number of reduction steps and, more importantly, to ensure that all $\Delta K$-nets associated with normalizing $\lambda K$-terms normalize. First, all core annihilations and erasures, including free variable erasure, are applied in any order. When no fan annihilations are possible, fan decays can be applied. Separately, when no replicator annihilations are possible, replicator decays can be applied. This ensures that fans and replicators are only eliminated through decay rules when they cannot be annihilated, which is more efficient. When no such reductions are possible (with the exception of eraser annihilation and free variable erasure, which can be applied at any time), core commutations are performed, augmented with lazy unpaired replicator merging. Commutations can only be applied when no annihilations, erasures, and decays are possible, otherwise, replications could be performed in subnets undergoing erasure, resulting, in the best case, in suboptimal behavior, and in the worst case in non-termination when termination is possible. Among the commutations, active pairs with unpaired replicators should be reduced last, as previously mentioned, in order to maximize the number of replicators that can be merged. When no commutations are possible, the aux fan replication rule is applied. Note that these are not separate phases, as in the $\Delta I$-Nets system, but rather a global partial ordering. Finally, when no further reduction operations can be applied, consecutive replicators can be merged since they are all unpaired, resulting in a canonical $\Delta K$-net.
Since all normal $\Delta$-nets are canonical, the $\Delta$-Nets systems are all confluent and possess the Church-Rosser property. The partial ordering of reduction operations specified for $\Delta K$-Nets is optimal in all $\Delta S$-Nets systems, $\forall S\in\Sigma$, omitting the appropriate reduction operations in each case. For ease of reference, the optimal $\Delta$-Nets partial order is summarized below. With the exception of reduction operations in group 0, which can be applied at any time, reduction operations in each numbered group can only be applied if no reduction operations in lower-numbered groups are possible. An asterisk indicates that a reduction operation is not a core interaction rule but a canonicalization rule instead. The systems to which each reduction operation applies are indicated in square brackets.
  1. 0.
    Unordered
    • $\bullet$
      Eraser annihilation [A,K]
    • $\bullet$
      Free-variable erasure [A,K]*
  2. 1.
    Deletions
    1. (a)
      Annihilations
      • $\bullet$
        Fan annihilation [L,A,I,K]
      • $\bullet$
        Replicator annihilation [I,K]
    2. (b)
      Erasures
      • $\bullet$
        Fan erasure [A,K]
      • $\bullet$
        Replicator erasure [K]
    3. (c)
      Decays
      Each should only be applied when no annihilations of the same agent type are possible, as annihilations delete agents in less interactions.
      • $\bullet$
        Fan decay [A,K]*
      • $\bullet$
        Complete replicator decay [K]*
  3. 2.
    Commutations
    • $\bullet$
      Fan replication [I,K]
    • $\bullet$
      Replicator replication [I,K]
    Before applying every commutation with an unpaired replicator:
    • $\bullet$
      it should be merged with all unpaired replicators which are connected to its auxiliary ports (unpaired replicator merging) [I,K]*
    • $\bullet$
      all auxiliary ports connected to erasers should be removed (partial unpaired replicator decay) [K]*
  4. 3.
    Expansions
    • $\bullet$
      Aux fan replication [I,K]*
  5. 4.
    Replicator Canonicalizations
    At this stage, only unpaired fan-in replicators remain, therefore all replicator trees are merged and (partially) decayed.
    • $\bullet$
      Unpaired replicator merging [I,K]*
    • $\bullet$
      Partial unpaired replicator decay [K]*
Definition.
For all $S\in\Sigma$ and for all $\lambda_{S}\in\Lambda_{S}$, if $\lambda_{S}$ is normalizing, $\lambda_{S}$ $\beta$-reduces to $\phi_{S}^{-1}(\Omega_{S}(\phi_{S}(\lambda_{S})))$ where $\Omega_{S}:\Delta_{S}^{p}\rightarrow\Delta_{S}^{c}$ reduces a proper $\Delta S$-net through interaction and canonicalization rules as defined in this section until it is normal and canonical.
Theorem.
Since $\phi_{S}^{-1}\circ\Omega_{S}$ is idempotent $\forall S\in\Sigma$, the set of all $\lambda S$-terms, $\Lambda_{S}$, is a projection of proper $\Delta S$-nets:
$\Lambda_{S}=\{\phi_{S}^{-1}(\Omega_{S}(\delta_{S}^{p}))\mid\forall\delta_{S}^{p}\in\Delta_{S}^{p}\},\ \forall S\in\Sigma$
Moreover, since $\Lambda_{S}$ is closed under $\beta$-reduction and $\Delta_{S}$ is closed under $\Delta_{S}$-interactions and canonicalizations, the $\lambda S$-calculus can be interpreted as a projection of $\Delta S$-Nets, $\forall S\in\Sigma$.
5.  Conclusion
The existence of non-linear non-canonical proper $\Delta$-nets reflects the additional degrees of freedom that gradual erasure and interior sharing introduce, which are not present in the $\lambda$-calculus. A given $\lambda A$-term, for example, can potentially be represented by many different proper $\Delta A$-nets, which differ only with respect to the progress of gradual erasure. Similarly, a given $\lambda I$-term can potentially be represented by many different proper $\Delta I$-nets, which differ only with respect to their sharing structure. As an example, take a proper $\lambda I$-net $n$, with $\phi_{I}^{-1}(\Omega_{I}(n))=MM$, for some $\lambda I$-term $M$. The $\lambda I$-net $n$ may have two distinct but equal subnets that each represent $M$, or it may have a single such subnet which is shared among the two occurrances in the application. In fact, the sharing structure of $n$ could be arbitrarily complex. These additional degrees of freedom allow $\Delta$-Nets to realize optimal reduction in the manner envisioned by Lévy, i.e., no reduction operation is performed which is rendered unneccessary later, and no reduction operation which is necessary is performed more than once.